Transform the vector A = 3i – 2j – 4k at P(2,3,3) to cylindrical coordinates.

ans.-3.6j – 4k


1
Expert's answer
2022-02-08T08:50:01-0500

Given:

"{\bf A} = 3\hat i – 2\hat j – 4\hat k"


The cylindrical coordinates are related to the cartesian ones by formulas

"\rho=\sqrt{x^2+y^2}\\
\phi=\tan^{-1}\frac{y}{x}\\
z=z"

In our case

"\rho=\sqrt{3^2+(-2)^2}=\sqrt{13}\\
\phi=\tan^{-1}\frac{-2}{3}=-34^\circ\\
z=-4"

Finally

"{\bf A}=(\sqrt{13},-34^\circ,-4)"